# Evaluation of uncertainty estimations for Gaussian process regression based machine learning interatomic potentials

### v1

@@ 3.2.2 Active learning via uncertainty sampling
Figure 5: Correlation of the GPR standard deviation and error across different datasets for GPR with Coulomb and GPR with SOAP. The samples were separated into equidistant bins by their uncertainty, and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical values from Eq. 15.
@@
3.3 Correlation analysis after applying active learning
It may be surprising that, in most cases, uncertainty sampling with the GPR standard deviation leads to a significantly worse MAE on the test set compared to random sampling. In Figure 5, we observe that predictions with the highest GPR standard deviation are having the highest bias. This would suggest that adding these configurations as training samples should be beneficial. The most extreme case is GPR with Coulomb matrices for aspirin, where the MAE after 800 iterations of uncertainty sampling with the GPR standard deviation as uncertainty is worse than for the initial model. To understand this behavior we show in Fig. 7 the correlation plots for Aspirin of a model trained via 800 iterations of uncertainty sampling with the GPR standard deviation as uncertainty estimation, as well as after iteratively adding 800 training configurations with the highest absolute error and adding 800 additional training configurations randomly. Similar to the model trained with 1000 random samples of Fig. 5, the initial model, which is trained with 200 random training configurations, shows an increasing negative bias for predictions with increasing GPR standard deviation, after GPR standard deviation exceeds a certain threshold. For predictions below this threshold, there is a range of rather low GPR standard deviation where the predictions have a positive bias. After adding 800 configurations randomly, this general form is still present. After uncertainty sampling, on the other side, the bias of predictions with a high standard deviation vanishes, while the positive bias, as well as the standard deviation of the errors, of predictions with low standard deviation, increases. A possible explanation for this is that, if we add the configuration for which the model predicts the highest GPR standard deviation in every iteration, we add the configuration which is most distant to the current training set. After many iterations, we have a training set with configurations which are maximally distant to each other. The training set now covers the borders of the configuration space, for which the initial model was showing a high bias, very well. However, predictions for some areas of the configuration space, which are less distant to the initial training configurations, are getting even worse, compared to the initial model with much less training configurations, when the training set overemphasizes the edges. In contrast to training via uncertainty sampling and adding training samples randomly, iteratively adding configurations for which the model has the highest absolute error results in a model which has a low error for predictions over the full range of different GPR standard deviations.
Figure 7: Correlation plots for GPR with Coulomb representations for Aspirin after 800 iterations of uncertainty sampling. The left plot in the first row shows the correlation plot with predicted GPR standard deviation vs. actual error for all configurations of the candidate pool of the initial GPR model that was trained with 200 random configurations. The right plot in the first row shows the uncertainty sampling results for 800 iterations. In the bottom row, correlation plots are shown for a GPR model trained with 800 iterations of uncertainty sampling with the GPR standard deviation as uncertainty estimation, as well as after iteratively adding 800 configurations with the highest absolute error and adding 800 additional training configurations randomly.
4 Discussion
The objective of this study was to evaluate different uncertainty estimators of GPR based MLIPs. To understand how uncertainty estimations correlate with the actual error, we plot both against each other and analyze if the errors are distributed according to the theoretical distributions, that result from the predictive distributions. For the GPR standard deviation, a clear correlation between an increasing uncertainty and increasing likelihood for larger errors is observed. For the two sets uncertainty, no correlation is observed, and the bootstrap uncertainty only shows a clear correlation in a few cases. However, even in cases where a correlation is observed, the actual error distributions often do not align with the error distributions that are predicted by the model. Therefore, the predictive distributions cannot quantify the error with valid confidence intervals. This also means it will not be possible to propagate uncertainty in a meaningful way. If the predictions of an MLIP are used to calculate further quantities, propagating uncertainty will not result in an accurate confidence intervals for the derived quantity. Still, if the uncertainty correlates with a higher likelihood for larger errors it provides a useful number, although a proper threshold might have to be found empirically. We found that specially predictions with higher uncertainties often will be highly overconfident, because the error is dominated by a high bias which is not captured quantitatively by the uncertainty. In contrast to our findings, Tran et al.38 found that the GPR standard deviation of their GPR model predicts underconfident confidence intervals for their exemplary dataset. The GAP model of Bartok et al.44 on the other hand predicts overconfident confidence intervals for their dataset of Ar dimers and trimers too, if the hyperparameters are optimized by maximizing the marginal likelihood, like we do. When optimizing the hyperparameters by optimizing the leave-one-out cross validation likelihood their GPR model predicts underconfident confidence intervals.
In ML theory, the error of a model is often separated into an irreducible part which results from the noise in the training data, as well as a bias and a variance part. We observe that predictions with a high uncertainty often have a high bias, which is not captured by the uncertainty in a quantitative way. Hüllermeier and Waegeman54 describe the GPR standard deviation as meaningful indicator for the total model uncertainty. However, if we have a model that has a bias in the form of systematical shifts of the predictive distributions in some areas of the input space it is not possible, that the standard deviation of a predictive distribution quantifies this shifts. Therefore, in general, a bias is associated with incorrect model assumptions, which makes it challenging to quantify the resulting error with the flawed model itself. For ensemble based uncertainties Heid et al.47 discussed that a bias can not be quantified.
@@ 4 Discussion
When sampling a fixed configuration space with uncertainty sampling, and the GPR standard deviation is used as uncertainty estimator, the test MAE does not improve compared to random sampling in cases where the predictive distributions aligned well with the error distributions, while the maximum absolute error and the test variance sometimes improve compared to random sampling. For cases where the predictions with increasing uncertainty have an increasing bias, which are most of the cases, the MAE is getting worse compared to random sampling. This aligns well with the findings of Uteva et al.9, where uncertainty sampling with the GPR variance performs worse than uncertainty sampling with the two sets uncertainty and adding samples with the highest absolute error, as well as training with maximum Latin hypercube data. This behaviour might be explained by the fact that the GPR standard deviation finds a maximally distant training set, which results in a model that overemphasizes the border of the configuration space, that was originally sampled in the fixed dataset. The borders of the configuration space, which is represented in the fixed dataset might be sampled more sparse while other areas are sampled more densely. A model that overemphasizes the border might perform slightly worse on samples of other areas of the configuration space, which are more frequent. While this means that uncertainty sampling can result in an increased average model error inside the fixed configuration space, it is likely to be beneficial for extrapolation tasks. This observation was also made by Podryabinkin et al.7 when they trained an MLIP with the most uncertain configurations of a fixed dataset, according to the D-optimality criterium. Compared to random sampling, the model shows a slightly increased root mean squared error and a slightly decreased maximum error, when testing on a test dataset, which was previously selected randomly from the fixed dataset, compared to random sampling, while MD simulations with this model are much more stable. The reduced data efficiency, when sampling a fixed configuration space with uncertainty sampling, suggest that for on-the-fly AL, where uncertainty estimations guide the exploration of configuration space, data efficiency may also be compromised. Contrary to this, Vandermause et al.3 were showing an increased data efficiency with their on-the-fly AL scheme applied to 5-element high entropy alloy structures.
Note that our tests were performed for energy predictions without forces. Other studies have reported that uncertainty estimations based on force predictions can be beneficial, because as forces are the derivative with respect to the atom coordinates, they carry more local information. Besides that, in general a higher model accuracy also leads to better uncertainty estimations. For the SOAP representations, parameters, which lead to higher resolution of the atomistic environments, would be possible. Still, we already use an atomistic kernel without sparsification, which in general leads to a much more accurate model than the classical GAP model with sparsification33.
In cases where the error distributions align well with the predictive GPR distributions, we find that data efficiency does not improve. If error distributions align with the predictive GPR distributions, even the errors of predictions with a high uncertainty are distributed around zero and many configurations for which the model has a high uncertainty actually have a low error. When we add samples with the actual highest error, we observe a significantly improved data efficiency. Therefore, if we would have a more linear correlation, i.e. the error is directly proportional to the uncertainty, between uncertainty estimations and errors instead of a correlation between uncertainty and how broad error distributions are distributed around zero, we might archive a much better data efficiency. One approach to archive this is to average over atomistic uncertainties55. If we would have calculated the bin averages of the absolute error in our correlation plots instead of the error, for cases where the predictive GPR distributions aligned well with the error distributions we would see a linear correlation between uncertainty and absolute error. However, this linear correlation is only archived because we average over many samples, for a single sample the most likely absolute error is zero. For atomistic models, like GPR with SOAP, the energy prediction can be seen as a sum of atomistic energy predictions of atomistic submodels. If forces are considered the force prediction for every atom can be seen as an atomistic submodel too. Now averaging the uncertainty prediction of these atomistic submodels can lead to a linear correlation between uncertainty estimation and absolute error.
@@ 5 Conclusion
In this study, we evaluated how different uncertainty estimations for GPR based MLIPs relate to the actual error and if sampling a fixed configuration space by iteratively choosing configuration with the maximum uncertainty can improve the model accuracy compared to random sampling. We find that ensemble based uncertainties often do not provide meaningful information about the error, while for the GPR standard deviation a clear correlation between increasing uncertainty and increasing likelihood for higher errors is observed. Still also the GPR standard deviation for most datasets is not able to quantify the prediction error correctly, therefore no accurate confidence intervals for model predictions can be derived. While an increasing GPR standard deviation correlates with an increasing bias in the model predictions, this bias is not captured quantitatively by the uncertainty estimation and often dominates the error of predictions with high uncertainty, resulting in highly overconfident confidence intervals. When sampling a fixed configuration space with configurations for which a GPR model predicts maximum GPR standard deviation, training configurations are found which are maximally distant to each other. This may lead to a more general model, that performs better for extrapolation tasks. Still, such a model can overemphasize the borders of the configuration space of a fixed dataset and perform on average worse on the fixed dataset than a model based on random sampling.
@@ Acknowledgment
The authors acknowledge support by the DFG through the DFG Priority Program SPP 2363 on “Utilization and Development of Machine Learning for Molecular Applications – Molecular Machine Learning” through the projects ZA 1175/4-1 and KL 1299/25-1. Some of the computations were carried out on the PLEIADES cluster at the University of Wuppertal, which was supported by the Deutsche Forschungsgemeinschaft (DFG, grant No. INST 218/78-1 FUGG) and the Bundesministerium für Bildung und Forschung (BMBF). The authors would also like to acknowledge the support of the ‘Interdisciplinary Center for Machine Learning and Data Analytics (IZMD)’ at the University of Wuppertal.
@@
To optimize the GPR hyperparameters we maximize the marginal log likelihood (MLL) using gpytorch45. We use the ADAM optimizer with a learning rate of $lr=0.1$ and 200 steps. Since prior assumption of the hyperparameters, i.e. the hyperparameter starting values in the optimization process, can significantly influence the outcome, we perform cross validation for all combinations of possible starting values shown in Table S2. In all cases we use 1000 random samples to optimize the hyperparameters, for which we perform 5-fold cross validation with 5 repetitions. In every fold we optimize the hyperparameters with the respective training samples and starting values. Then we choose the hyperparamter starting values which resulted in the lowest loss on average over all repetitions and folds and use them as starting values for maximizing the MLL with all 1000 random samples.
@@
Table S1: Possible starting values for hyperparameters for the maximization of the MLL.
@@
Figure S1: Correlation plots of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset. On the one hand the predictions are the mean of the submodel predictions of the two sets or bootstrap ensemble, on the other hand the predictions are from a single GPR model, trained on the full training set, from which subsets are used to train the ensemble submodels.
@@
Figure S2: Correlation plots of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for aspirin from the rMD17 dataset. On the one hand the predictions are the mean of the submodel predictions of the two sets or bootstrap ensemble, on the other hand the predictions are from a single GPR model, trained on the full training set, from which subsets are used to train the ensemble submodels.
@@
S4 Comparison of Uncertainty error correlation between several molecules
@@
Figure S3: Correlation of the two sets uncertainty and error across different datasets for GPR with Coulomb and GPR with SOAP. The samples were separated into equidistant bins by their uncertainty and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical error distributions. Note that for porphyrin and GPR with Coulomb the axis ranges and and bin width are adapted in comparison to the correlation plot for the GPR standard deviation for better visibility.
@@
Figure S4: Correlation of the bootstrap uncertainty and error across different datasets for GPR with Coulomb and GPR with SOAP. The samples were separated into equidistant bins by their uncertainty and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical error distributions. Note that for porphyrin and GPR with Coulomb the axis ranges and and bin width are adapted in comparison to the correlation plot for the GPR standard deviation for better visibility.
S5 Comparison of uncertainty sampling between several molecules
(a) GPR with Coulomb
(b) GPR with SOAP
Figure S5: Maximum absolute error of every uncertainty sampling iteration for GPR with Coulomb and SOAP representations for different datasets. We randomly selected 200 samples as initial training samples and 2000 samples as test samples. From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample. Additionally we randomly added a new sample in every iteration and we added the sample with the highest actual absolute error.
(a) GPR with Coulomb
(b) GPR with SOAP
Figure S6: Test variance of every uncertainty sampling iteration for GPR with Coulomb and SOAP representations for different datasets. We randomly selected 200 samples as initial training samples and 2000 samples as test samples. From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample. Additionally we randomly added a new sample in every iteration and we added the sample with the highest actual absolute error.

### v2

@@ 3.2 Local calibration analysis reveals model bias
Figure 4: Calibration curves and extended reliability diagrams of the GPR standard deviation of GPR with Coulomb applied to different datasets. For the extended reliability diagrams, samples were separated into equidistant bins by their uncertainty, and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical values from Eq. 14.
3.3 Uncertainty sampling results in worse error than random sampling
In Fig. 5, the MAE for every iteration of uncertainty sampling across all molecules is shown. In most cases, uncertainty sampling with the GPR standard deviation results in a worse MAE compared to adding samples randomly. Only for benzene and porphyrin with Coulomb matrices, uncertainty sampling results in a similar MAE compared to adding samples randomly. Note that in the local calibration analysis, this were exactly the cases where the extended reliability diagrams did not show a significant bias for predictions with high uncertainties. The two sets uncertainty performs very similar to random sampling, as expected, since the extended reliability diagrams demonstrate that it does not provide any meaningful information about the error. Interestingly, uncertainty sampling with the bootstrap aggregation uncertainty results in a worse MAE than random sampling in almost all cases. When samples corresponding to predictions with the highest absolute error were added to the training set, the MAE improved significantly for all cases, showing that in principle there is a relevant potential for an improved data efficiency through better sampling strategies.
Our findings align well with those of Uteva et al.[9], who reported that uncertainty sampling with the GPR variance performs worse than uncertainty sampling with the two sets uncertainty, adding samples corresponding to predictions with the highest absolute error, and training with maximum Latin hypercube data. The reduced data efficiency observed when sampling a fixed configuration space using uncertainty sampling suggests that, in on-the-fly active learning (AL), where uncertainty estimations guide the exploration of the configuration space, data efficiency may also be compromised. In contrast, Vandermause et al.[3] demonstrated increased data efficiency with their on-the-fly AL scheme applied to five-element high-entropy alloy structures.
Even if the MAE is not improved by uncertainty sampling, a model trained via uncertainty sampling may perform more consistently across different data points in the test set. To analyze this, we present results for the maximum absolute test error and the test variance in Figs. S9 and S10 of the supplementary material. Indeed, the maximum absolute error of the test set improves for uncertainty sampling with the GPR standard deviation in all cases, except for aspirin. The test variance, however, is similar to that of random sampling in most cases or deteriorates, with improvements observed only for benzene and O-HBDI with Coulomb. When training samples corresponding to predictions with the highest absolute error were added, both metrics improved significantly for all cases.
@@ 3.3 Uncertainty sampling results in worse error than random sampling
3.4 More general model after uncertainty sampling?
The significantly worse MAE for uncertainty sampling with the GPR standard deviation compared to random sampling might seem surprising, as we observe a correlation between increasing GPR standard deviation and increasing bias. This suggests that adding configurations corresponding to the highest GPR standard deviation as training samples should be beneficial. To analyze this behavior, in Fig. 6 we show extended reliability diagrams after uncertainty sampling with the GPR standard deviation and after random sampling for GPR with Coulomb applied to Aspirin and SMA. For SMA, we show the plots for the models after all 4800 iterations, while for Aspirin, the plots are shown after only 800 iterations, because at that point, the model trained via uncertainty sampling with the GPR standard deviation performs even worse than the initial model. Additionally, we show the plot for the initial GPR model trained with 200 random samples. In both cases, it can be observed that the initial model not only has a high bias for predictions with high uncertainties but also a smaller bias in the opposite direction for predictions with very low uncertainty. After uncertainty sampling, the high bias for predictions with high uncertainty is eliminated in both cases. However, the bias for predictions with low uncertainty is higher than after random sampling and, for Aspirin, even higher than for the initial model. The standard deviation of errors of the respective bins with low uncertainty is higher than after random sampling too.
@@ 3.4 More general model after uncertainty sampling?
These observations could be explained by the fact that uncertainty sampling with the GPR standard deviation finds a maximally distant training set. This may result in a model that strongly emphasizes the borders of the configuration space originally sampled in the fixed dataset. In such fixed datasets, the borders of the configuration space might be sampled more sparsely, while other areas are sampled more densely. A model that emphasizes the borders strongly might perform slightly worse on samples from other areas of the configuration space, which are more frequent. While this means that uncertainty sampling can result in an increased average model error within the fixed configuration space, it is likely to be beneficial for extrapolation tasks. This observation was also made by Podryabinkin et al.[7], who trained an MLIP using the most uncertain configurations of a fixed dataset, based on the D-optimality criterion. Compared to random sampling, the model shows a slightly increased root mean squared error and a slightly decreased maximum error when tested on a dataset randomly selected from the fixed dataset. However, MD simulations using this model were significantly more stable.
(a) Aspirin
(b) SMA
Figure 6: Extended reliability diagrams of GPR with Coulomb representations for aspirin and SMA are presented. These plots compare the initial GPR models trained on 200 random configurations with those obtained after uncertainty sampling using the GPR standard deviation, as well as models updated by adding samples randomly. For better visibility, the plots for SMA use different bin sizes and axis limits between the initial model and those following uncertainty sampling. Note that the axis values also differ from those in Fig. 4.
@@ 4 Conclusion
In this study, we evaluated calibration of different uncertainty measures for GPR based MLIPs and their performance when applied to active learning, where we iteratively choose the configuration with the maximum uncertainty from a fixed configuration space. Regarding calibration, it becomes evident how important it is to consider not only global calibration, which averages over the entire test dataset, but also local calibration, which considers specific groups of predictions. Ensemble-based uncertainty measures appear to be a poor choice for GPR-based MLIPs, as they already show poor global calibration. In contrast, the GPR standard deviation demonstrates relatively good global calibration. However, when calibration is evaluated locally, the GPR standard deviation can show very poor calibration. We find that, in particular, groups of predictions with a high GPR standard deviation often have a high bias that is not captured quantitatively by the uncertainty estimation. Therefore, without further knowledge, the GPR standard deviation should not be interpreted as a quantitative measure of potential error ranges, such as confidence intervals derived from the GPR standard deviation. Still, we find that an increasing GPR standard deviation correlates with an increasing bias and can therefore identify predictions with high bias and high error. When sampling a fixed configuration space by selecting configurations for which a GPR model predicts the maximum GPR standard deviation, the resulting training configurations are maximally distant from each other. This may lead to a more general model, that performs better in extrapolation tasks. However, such a model may overemphasize the borders of the configuration space in a fixed dataset, resulting in worse average performance on the fixed dataset compared to a model trained using random sampling.
@@ Acknowledgment
The authors acknowledge support by the DFG through the DFG Priority Program SPP 2363 on “Utilization and Development of Machine Learning for Molecular Applications – Molecular Machine Learning” through the projects ZA 1175/4-1 and KL 1299/25-1. Some of the computations were carried out on the PLEIADES cluster at the University of Wuppertal, which was supported by the Deutsche Forschungsgemeinschaft (DFG, grant No. INST 218/78-1 FUGG) and the Bundesministerium für Bildung und Forschung (BMBF). The authors would also like to acknowledge the support of the ‘Interdisciplinary Center for Machine Learning and Data Analytics (IZMD)’ at the University of Wuppertal. The authors also acknowledge the use of ChatGPT-4o, for assistance with language refinement, grammar corrections, and reformulations in the preparation of this manuscript.
Data availability statement
The data that support the findings of this study are openly available at the following URL: xxx
@@
To optimize the GPR hyperparameters we maximize the marginal log likelihood (MLL) using GPyTorch[47]. We use the ADAM optimizer with a learning rate of $lr=0.1$ and 200 steps. Since the initial guess of the hyperparameters in the optimization process can significantly influence the outcome, we perform cross validation for all combinations of possible initial guesses shown in Table S1. In all cases, we use 1000 random samples to optimize the hyperparameters, for which we perform 5-fold cross validation with 5 repetitions. In every fold, we optimize the hyperparameters with the respective training samples and starting values. Then we, choose the hyperparameter initial guesses, which resulted in the lowest loss on average over all repetitions and folds, and use them as intitial guesses for maximizing the MLL with all 1000 random samples.
@@
Table S1: Possible initial guesses for hyperparameters for the maximization of the MLL.
@@
Figure S1: Extended reliability diagrams of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset.
@@
Figure S2: Extended reliability diagrams of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for aspirin from the rMD17 dataset.
@@
S4 Calibration analysis
Figure S3: Calibration curves and extended reliability diagrams of the two sets uncertainty of GPR with Coulomb applied to different datasets.
Figure S4: Calibration curves and extended reliability diagrams of the bootstrap uncertainty of GPR with Coulomb applied to different datasets.
Figure S5: Calibration curves and extended reliability diagrams of the GPR standard deviation of GPR with SOAP applied to different datasets.
Figure S6: Calibration curves and extended reliability diagrams of the two sets uncertainty of GPR with SOAP applied to different datasets.
Figure S7: Calibration curves and extended reliability diagrams of the bootstrap uncertainty of GPR with SOAP applied to different datasets.
S5 Explicit error distributions of separate bins
(a) Aspirin
(b) O-HBDI
Figure S8: Predicted error distribution and distribution of true errors for selected bins of the reliability diagram of GPR with SOAP for Aspirin and O-HBDI.
S6 Other metrics uncertainty sampling
@@
Figure S9: Maximum absolut error of uncertainty sampling for GPR with Coulomb and SOAP representations for different datasets.
@@
Figure S10: Test variance of uncertainty sampling for GPR with Coulomb and SOAP representations for different datasets.